OpenStudy (anonymous):

Which of the following ordered pairs is a solution to the inequality y >= 4x - 2? (0, -4) (0, 0) (0, -3) (0, -5)

OpenStudy (anonymous):

(0,0) 0>-2

Parth (parthkohli):

Plug in the numbers and see what satisfies.

OpenStudy (anonymous):

0 >= 4*-4 -2?

Parth (parthkohli):

Yeah... just keep plugging like that and see if any of the inequalities satisfy.

OpenStudy (anonymous):

0 >+ 4*0 - 2

OpenStudy (anonymous):

0 >= 4*-3 - 2

Parth (parthkohli):

@Solmyr the solution sets are written in the form \((x,y)\).

OpenStudy (anonymous):

0 >= 4*-5 -2

OpenStudy (anonymous):

so i need to put the 0 as X

Parth (parthkohli):

Instead, make them look like these:\[-4\ge 4(0) - 2 \\ 0 \ge 4(0) - 2\\ -3 \ge 4(0) - 2\\-5 \ge 4(0) - 2 \]

OpenStudy (anonymous):

ok srry did realize that

OpenStudy (anonymous):

didnt

Parth (parthkohli):

All good :)

OpenStudy (anonymous):

thank you i understand it now

Parth (parthkohli):

You're welcome :)